f(x)=x^2 when x>=0 and f(x)=-x^2 when x <0. find 2f(x)>=f(4-x^2)
Answer #1

Ques solution x2 N 7,0 Given Yolution flm) -12 nco then. N TO 2012 2 fon) -2 n2 i aco graph of 2 fm) will be an follows 1 y-aNow deduce from f 14-32) given function for we can an, 4-112 to -☺ f (4-2) = (4-02)2 -(4-22) 2 ; 4-12 co Now, 4.227,0 -) 212<How to find intersection point of 2 fois and 2 (4-22) 2 (4-2232 -) ㅕ tun?) 222 = (4-72) for OLN<2 and -2012 for n<-2 2012 = 2